# Proto-CLIP: Vision-Language Prototypical Network for Few-Shot Learning

### v1

@@ 4.1 Ablation Studies
Adapter Types and Learnable Text Memory. Since the 12 datasets have different characteristics, we found that varying adapter types and whether to learn the text memory or not affect performance. Table 2 summarizes the result of this ablation study. The architectures of the MLP-based adapter and the convolution-based adapter are illustrated in Fig. 3. “2xConv” indicates using 2 convolution layers as shown in Fig. 3, while “3xConv” uses 3 convolution layers in the adapter where we add a $32@3\times 3\times 32$ convolution layer in the middle. By checking the best accuracy for each dataset, we can see that there is no consensus on which adapter and trainable text memory to use among these datasets. Therefore, we select the best configuration on the adapter and learnable text memory for each dataset in the following experiments. Learning both image memory and text memory can obtain aligned image-text prototypes. Fig. 4 visualizes the image-text prototypes in the FewSOL dataset [15] before and after training.
@@
Table 2: Ablation study of query adapters with $K=16$ and Proto-CLIP-$F$. In all cases, the adapter and the visual memory keys are trained. In case of tie, an underlined setup was used.
Loss functions. We have introduced three different loss functions in Sec. 3: $\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3}$. We analyze the effects of these loss functions in Table 3. We can see that i) the $\mathcal{L}_{1}$ loss function is essential since it drives the classification of the query images; ii) Overall, both $\mathcal{L}_{2}$ and $\mathcal{L}_{3}$ loss functions for prototype alignment contribute to the performance, which verifies our motivation of aligning image and text prototypes for few-shot classification.
@@
Table 3: Loss function ablation results with the ResNet50 backbone and shot $K=16$.
Backbones. Table 4 shows the results of using different backbone networks on the FewSOL dataset [15]. In general, better backbones can learn more powerful feature representations and consequently improve the classification accuracy. CLIP vision transformer backbones achieve better performance than CLIP ResNet backbones.
Model Adapter TextM Backbone
@@
Table 4: Backbone ablation study. Dataset=FewSOL-52 [15]. $K=16$. Model=Proto-CLIP-$F$. “TextM” indicates whether to train text memory.
4.2 Comparison with Other Methods
@@
Table 5 shows the performance of Proto-CLIP compared to the state-of-the-art methods using CLIP for few-shot learning in the literature: Linear-Probe CLIP [1], CoOp [8], CLIP-Adapter [9] and Tip-Adapter [8]. We follow these methods and use ResNet50 backbone for this comparison. The fine-tuned variant of Tip-Adapter “Tip-F” is the most competitive method compared to ours. The performance of Proto-CLIP on very few shots, i.e., 1 shot and 2 shots is inferior compared to Tip-F. When the number of shots increases to 4, 8 and 16, the fine-tuned variants of Proto-CLIP outperform Tip-F. Proto-CLIP-$F$-$Q^{T}$ performs better than Proto-CLIP-$F$ on most datasets by using the data augmentation of query images during training.
# classes 1,000 100 37 196 10 100 397 47 102 101 101 52
1 shots
CLIP-A [9] 61.20 17.49 85.99 55.13 61.40 88.60 61.30 45.80 73.49 76.82 62.20 -
2 shots
CLIP-A [9] 61.52 20.10 86.73 58.74 63.90 89.37 63.29 51.48 81.61 77.22 67.12 -
4 shots
CLIP-A [9] 61.84 22.59 87.46 62.45 73.38 89.98 65.96 56.86 87.17 77.92 69.05 -
8 shots
CLIP-A [9] 62.68 26.25 87.65 67.89 77.93 91.40 67.50 61.00 91.72 78.04 73.30 -
16 shots
CLIP-A [9] 63.59 32.10 87.84 74.01 84.43 92.49 69.55 65.96 93.90 78.25 76.76 -
Table 5: Few-shot classification results on different datasets using the ResNet50 backbone.
4.3 Real World Experiments
@@
As an application, we have built a robotic system to verify the effectiveness of Proto-CLIP for object recognition in the real world. Fig. 5 illustrates our pipeline for the system. It takes human instruction in the form of voice commands as input such as “pick something” or “grasp something”. The system first applies Automatic Speech Recognition (ASR) to convert voice input to text using OpenAI Whisper [33]. Then the system grounds the noun in the human instruction into a target object observed from an input image. This is achieved by joint object segmentation and classification. We utilize unseen object instance segmentation [34] to segment objects in cluttered scenes and then classify each segmented object with Proto-CLIP. By matching the noun with the class labels, the system can ground the target in the image. Once the target object is recognized, we use Contact-GraspNet [35] for grasp planning and MoveIt motion planning toolbox [36] to pick and place the target. See the supplementary material for more real-world results.
5 Limitations
Proto-CLIP performs poorly in low-shot regimes, as is evident from Table 5. A hyperparameter grid search is necessary for each new dataset, following the methodology of Tip-Adapter. This requirement applies to every combination of the new dataset and the backbone. Embracing the diversity of datasets, our system thrives on the need for different set-ups. When encountering a new dataset, we actively compare the effectiveness of $F$ and $F$-$Q^{T}$ to determine the optimal choice. This dynamic approach transforms the potential weakness into a strength, allowing us to adapt and maximize performance for every unique dataset. During our observations, we discovered that data transformations play a crucial role in building the cache model.
6 Conclusion and Future Work
@@
We have introduced a novel method for few-shot learning based on the CLIP vision-language model. Our method learns image prototypes and text prototypes from few-shot training examples and aligns the corresponding image-text prototypes for classification. The model is equipped with learnable image memory and text memory for support images and a learnable adapter for query images. Compared to previous CLIP-based few-shot learning methods, our method is flexible in configuring these learnable components, resulting in powerful learned models.
Good feature representation is the key in few-shot learning. Future work includes how to further improve feature representation learning compared to CLIP models. One idea is to adapt more powerful vision-language models such as GPT variants. The FewSOL dataset also provides multiview and depth information about objects. Exploring this 3D information in few-shot object recognition is also a promising direction.
Acknowledgments

### v3

@@ IV-A Ablation Studies
Adapter Types and Learnable Text Memory. Since the 12 datasets have different characteristics, we found that varying adapter types and whether to learn the text memory or not affect performance. Table III summarizes the result of this ablation study. Visual data plays a crucial role in image recognition when compared to textual information. Therefore, visual memory keys are consistently trained, regardless of the circumstances. The architectures of the MLP-based adapter and the convolution-based adapter are illustrated in Fig. 3. “2xConv” indicates using 2 convolution layers as shown in Fig. 3, while “3xConv” uses 3 convolution layers in the adapter where we add a $32@3\times 3\times 32$ convolution layer in the middle. By checking the best accuracy for each dataset, we can observe that there is no consensus on which adapter and trainable text memory setup to use among these datasets. Therefore, we select the best configuration on the adapter and learnable text memory for each dataset in the following experiments. Learning both image memory and text memory can help to yield aligned image-text prototypes. Fig. 4 visualizes the image-text prototypes in the FewSOL dataset [15] before and after training. For Proto-CLIP-$F$, unless specified otherwise, both the adapter and the visual memory keys are trained in all scenarios.
# classes 1000 100 37 196 10 100 397 47 102 101 101 52
1 shot
CLIP-Adapter [9] 61.20 17.49 85.99 55.13 61.40 88.60 61.30 45.80 73.49 76.82 62.20 -
2 shots
CLIP-Adapter [9] 61.52 20.10 86.73 58.74 63.90 89.37 63.29 51.48 81.61 77.22 67.12 -
4 shots
CLIP-Adapter [9] 61.84 22.59 87.46 62.45 73.38 89.98 65.96 56.86 87.17 77.92 69.05 -
8 shots
CLIP-Adapter [9] 62.68 26.25 87.65 67.89 77.93 91.40 67.50 61.00 91.72 78.04 73.30 -
16 shots
CLIP-Adapter [9] 63.59 32.10 87.84 74.01 84.43 92.49 69.55 65.96 93.90 78.25 76.76 -
TABLE II: Few-shot classification results of various CLIP based few shot learning methods on different datasets across various shots using the CLIP ResNet50 backbone.
Loss functions. We have introduced three different loss functions in Sec. III: $\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3}$. We analyze the effects of these loss functions in Table IV. We can see that i) the $\mathcal{L}_{1}$ loss function is essential since it drives the classification of the query images; ii) Overall, both $\mathcal{L}_{2}$ and $\mathcal{L}_{3}$ loss functions for prototype alignment contribute to the performance, which verifies our motivation of aligning image and text prototypes for few-shot classification.
@@
TABLE III: Results of ablation study of various query adapter types and textual memory bank training using the CLIP ResNet50 backbone with $K=16$ on Proto-CLIP-$F$. In case of a tie, the underlined setup was selected randomly.
@@
TABLE IV: Ablation study of various Loss functions using the CLIP ResNet50 backbone and $K=16$. The best performing model architectures for each dataset from Table III are used here.
Backbones. Table V shows the results of using different backbone networks on the FewSOL dataset [15]. In general, better backbones can learn more powerful feature representations and consequently improve the classification accuracy. CLIP vision transformer backbones achieve better performance than CLIP ResNet backbones.
Model Adapter TTM Backbone
@@
TABLE V: Backbone ablation study. Dataset=FewSOL-52 [15]. $K=16$. TTM=‘Train-Text-Memory’.
Dataset Method 1 2 4 8 16 32 64
ImageNet [5] Tip [10] 60.70 60.96 60.98 61.45 62.01 62.51 62.88
Proto-CLIP 60.31 60.64 61.30 62.12 62.77 62.98 63.23
Tip-F [10] 61.13 61.69 62.52 64.00 65.51 66.58 67.96
Proto-CLIP-$F$ 60.32 60.64 61.30 63.92 65.75 66.47 65.36
Proto-CLIP-$F$-$Q^{T}$ 59.12 60.48 61.80 64.03 65.91 66.71 66.90
FewSOL-52 [15] Tip [10] 27.30 26.22 28.70 29.22 28.87
Proto-CLIP 27.09 28.35 29.13 29.83 29.96
Tip-F [10] 27.91 27.43 29.13 32.43 34.04
Proto-CLIP-$F$ 22.22 26.17 27.09 33.26 35.22
Proto-CLIP-$F$-$Q^{T}$ 21.65 25.91 30.30 32.70 34.70
TABLE VI: Shots ablation results. Backbone=‘CLIP ResNet50’.
Shots. Table VI displays the results of using different numbers of shots on ImageNet [5] and FewSOL [15]. With more shots for training, the classification accuracy is improved accordingly. The choice of K=16 for our experiments aligns with the prevalent practice in the field of vision-language few-shot learning. This specific value has been widely adopted, as evidenced in various scholarly works such as  [9, 8, 10] Moreover, given our specific emphasis on the few-shot context, it appeared prudent to exercise caution when surpassing a particular threshold, specifically 16 in our case. As a result, we embarked on an ablation study involving the ImageNet [5] dataset. This particular dataset holds the largest number of classes (1000) and thus provided a suitable platform for investigating shots values beyond 16, such as 32 and 64. Despite our intention to explore 128 shots, our experimental hardware’s memory limitations prohibited us from pursuing this avenue. Additionally, FewSOL is valuable for few-shot object learning, especially in robotics. We capped shots at 16 for FewSOL as average number of samples per class in FewSOL hovers around 15. Consequently, we conjectured that going beyond might yield diminishing learning returns. These insights are detailed in Table VI.
IV-B Comparison with Other Methods
@@
Table II shows the performance of Proto-CLIP compared to the state-of-the-art methods using CLIP for few-shot learning in the literature: Linear-Probe CLIP [1], CoOp [8], CLIP-Adapter [9] and Tip-Adapter [8]. We follow these methods and use CLIP’s ResNet50 backbone for this comparison. The fine-tuned variant of Tip-Adapter “Tip-F” is the most competitive method compared to ours. The performance of Proto-CLIP on very few shots, i.e., 1 shot and 2 shots is inferior compared to Tip-F. When the number of shots increases to 4, 8 and 16, the fine-tuned variants of Proto-CLIP outperform Tip-F. The enhanced performance of our proposed Proto-CLIP method can be attributed to its reliance on robust image and textual prototypes, which subsequently leads to improved classification accuracy. Therefore, our model benefits from more than 4 shots, while it is not as good as Tip-F when using 1 shot and 2 shots. Proto-CLIP-$F$-$Q^{T}$ performs better than Proto-CLIP-$F$ on most datasets by using the data augmentation of query images during training [^2] For more details, please see the supplementary material on the project page..
IV-C Real World Experiments
@@
As an application, we have built a robotic system to verify the effectiveness of Proto-CLIP for object recognition in the real world. Fig. 5 illustrates our pipeline for the system. It takes human instruction in the form of voice commands as input such as “pick something” or “grasp something”. The system first applies Automatic Speech Recognition (ASR) to convert voice input to text using OpenAI Whisper [35]. Then the system grounds the noun in the human instruction into a target object observed from an input image. This is achieved by joint object segmentation and classification. We utilize unseen object instance segmentation [36] to segment objects in cluttered scenes and then classify each segmented object with Proto-CLIP. By matching the noun with the class labels, the system can ground the target in the image. Once the target object is recognized, we use Contact-GraspNet [37] for grasp planning and MoveIt motion planning toolbox [38] to pick and place the target2.
V CONCLUSIONS
@@
We have introduced a novel method for few-shot learning based on the CLIP [1] vision-language model. Our method learns image prototypes and text prototypes from few-shot training examples and aligns the corresponding image-text prototypes for classification. The model is equipped with learnable image memory and text memory for support images and a learnable adapter for query images. Compared to previous CLIP-based few-shot learning methods, our method is flexible in configuring these learnable components, resulting in powerful learned models. Good feature representation is the key in few-shot learning. Future work includes how to further improve feature representation learning compared to CLIP models. One idea is to adapt more powerful vision-language models such as GPT variants. The FewSOL [15] dataset also provides multiview and depth information about objects. Exploring this 3D information in few-shot object recognition is also a promising direction.
ACKNOWLEDGMENT